i need to solve for all real roots
(sqrt)(2x^2+x-12)=x

Answer by venugopalramana(3286)

i need to solve for all real roots
(sqrt)(2x^2+x-12)=x............squaring..
2x^2+x-12=x^2
x^2+x-12=0
x^2+4x-3x-12=0
x(x+4)-3(x+4)=0
(x-3)(x+4)=0
x=3....or....x=-4